# Broken symmetry solutions in one-dimensional lattice models via many-body perturbation theory

### v1

@@ V Long-range interaction
We note in passing that different parametrizations have also been employed in similar models accounting for a non-local electron-electron repulsion Ohno (1964); Kaasbjerg and Thygesen (2010); Strange and Thygesen (2012). All these parametrizations, however, share the same $V_{ij}\propto 1/R_{ij}$ behavior at large distances. The model depends on two dimensionless parameters: the ratio $U/t$ defines the strength of the correlation, while the ratio $V/U$ sets the strength of the non-local interaction term with respect to the local one. We take $V/U\leq 1$ as we expect inter-site electron repulsion to be always smaller than the on-site one.
@@ V.1 Results
The nature of the exact ground state can be further characterized by the introduction of appropriate correlation functions involving two density operators. In particular, even if the exact ground state shares the same symmetry of the non-interacting one, it can still develop an AFM or CDW character, which can be measured by the evaluation of certain correlation functions. This is, for instance, the case of the Hubbard model, whose ground state develops strong AFM correlations in the large $U/t$ regime without breaking explicitly the symmetry of the Hamiltonian Joost et al. (2022); Honet et al. (2022).
@@
which measures the overall charge dimerization on two adjacent sites. The factor of $1/4$ in Eq. (23) takes into account the number of all spin combinations entering the sum and makes Eq. (23) comparable to Eq. (22). The smaller the value of the correlation functions in Eqs. (22) and (23), the larger the corresponding AFM or CDW character. While these correlation functions are easily accessible in the exact calculations, as they can be directly computed through the action of field operators on the many-body wavefunctions, they can not be directly evaluated in MBPT calculations, where we have mostly access to the one-particle Green’s function ($C^{AFM}$ in Eq. (22) is easily accessible only in the Hubbard model, where the local nature of the electron-electron interaction allows one to extract this correlation function from the knowledge of the total energy and interaction energy Joost et al. (2022); Honet et al. (2022)). For the broken symmetry solutions obtained within approximate MBPT calculations, we then consider a “factorized” version of the former correlation functions:
@@
Importantly, for this particular system the symmetry breaking at the MBPT level, while unphysical, follows from the development of analogous correlations in the exact system. Interpreting the ratio $V/U$ in terms of the distance between the sites of the chain – i.e. a small value of $V/U$ representing a large inter-site distance we can see that AFM correlations dominate in the large distance regime as observed for hydrogen chains Motta et al. (2020), while CDW correlations are relevant at smaller inter-site distances Barborini et al. (2022).
@@ V.2 Solutions of the SSE
It is interesting to stress that, due to the simplified nature of the lattice model, the symmetric solution leads to a uniform charge density, which is then the same for all the values of the model parameters. This means that, for the symmetric solutions, the KS potential $v_{KS}$ is just a constant term that aligns the highest occupied molecular (quasiparticle) orbital (HOMO) of the non-interacting ground state to the HOMO of the interacting symmetric state. In particular, in the case of chains having a metallic non-interacting ground state (see Supplemental Material sup ), the Kohn-Sham Green’s function $G_{KS}$ describes a metallic system, while the system displays a gap at the MBPT level. We stress here that in order to converge the GW calculations in the case of the symmetric state, the self-consistency loop is initialized with a gapped GF.
Instead, in the case of broken symmetry solutions, $G_{KS}$ describes a gapped system at finite values of $U$ and $V$. The ratio between the derivative discontinuity $\Delta_{xc}$ and the quasiparticle energy gap is reported in Fig. 8.
@@
We observe that the derivative discontinuity for the $N=16$ chain is large and accounts for $70\%$ to $80\%$ of the quasiparticle energy gap. Similar results have also been observed in model 1D semiconductors with long-range electronic interactions Schonhammer and Gunnarsson (1987). As an interesting point, the ratio between $\Delta_{xc}$ and $E_{g}^{QP}$ gets smaller by increasing the parameter $V/U$ for the CDW solution, while the opposite happens for the AFM solution. For the $N=16$ chain the value of $V/U$ at which the two curves of Fig. 8 cross is also the value of $V/U$ at which the energy of the AFM solution becomes larger than that of the CDW one at the level of GW. Having a large derivative discontinuity, the energy gap of $G_{KS}$ in the case of broken symmetry solutions is severely underestimated with respect to the quasiparticle gap.
In Fig. 9 we report the DOS of the periodic $N=16$ chain at $U/t=4$ and $V/U=0.75$. The DOS has been evaluated in the case of the GW for the symmetric as well as broken symmetry solutions. The solid line reports the DOS of the self-consistent GF, while the dashed line reports the DOS of $G_{KS}$. The position of the HOMO level is also highlighted. As discussed in Sec. III.2, the system of linear equations defined by the SSE, Eq. (II.2), is solved requiring that the HOMO level of $G_{KS}$ is aligned to that of $G$.
Figure 9: DOS of the $N=16$ chain with periodic boundary conditions at $U/t=4$ and $V/U=0.5$ for a long-range Coulomb interaction. We report the DOS for all states with different symmetry. The red line is related to the GW DOS, while the green line is the DOS of the corresponding non-interacting $G^{KS}$. The HOMO quasiparticle level is marked with a downwards triangle and shifted to $\omega=0$ in all cases. The frequency is scaled with the hopping parameter $t$ and the DOS is reported in arbitrary units.
@@ VI Conclusions
In this work we study one-dimensional lattice models with an approximate many-body perturbation theory (MBPT) treatment of the electron-electron interaction. The Dyson equation is solved in a fully self-consistent manner via the algorithmic inversion method based on the sum-over-pole representation of dynamical operators (AIM-SOP) Chiarotti et al. (2022); Chiarotti (2023); Chiarotti et al. (2024). The excellent agreement between thermodynamic and spectral properties of our work with the results of Refs. Puig von Friesen et al. (2010); Di Sabatino et al. (2021) shows that the AIM-SOP formulation is capable of high numerical accuracy. In particular, the fully self-consistent implementation allows one to study in detail the total energy of broken symmetry solutions by evaluating unambiguously the value of the Klein functional at self-consistency. Notably, in this work the Klein functional has been evaluated in an analytical manner following Ref. Ferretti et al. (2024).
Moving to symmetry-broken magnetic solutions, we obtain an excellent agreement with the results of Ref. Joost et al. (2022), regarding antiferromagnetic solutions in the 1D Hubbard model studied within the second Born approximation. As already discussed in Refs. Joost et al. (2022); Honet et al. (2022), the MBPT treatment of the interaction may result in unphysical solutions breaking the symmetry of the Hamiltonian and in disagreement with the symmetry of the exact solution Lieb and Wu (1968, 2003b). This is the case of antiferromagnetic solutions that are found in the 1D Hubbard model at large values of the interaction strength $U/t$.
@@
The role of broken symmetry solutions is then investigated in a model with long-range interactions in one-dimensional chains with periodic boundary conditions. In this case, while the exact ground state has again a homogeneous charge density, broken symmetry solutions with a charge density wave (CDW) or antiferromagnetic (AFM) ordering can be found within an approximate MBPT treatment. We show the example of a $N=4$ chain where the unphysical MBPT broken symmetry solutions can be traced back to the development of an AFM or CDW character in the exact ground state. This character can be measured in the exact solution by evaluating correlation functions involving two density operators.
Remarkably the broken symmetry solutions found within GW reproduce the trend of the two-density correlation functions, signaling that the symmetry breaking experienced in the approximate MBPT calculation follows from the development of analogous correlations present also in the exact system. In particular, the CDW character gets enhanced when the strength of the intersite electronic repulsion $V$ approaches the strength of the local electronic repulsion $U$. Similarly to what observed in the $N=4$ chain, broken symmetry solutions in one-dimensional periodic chains are found when the non-interacting ground state of our model, Eq. (20), is metallic: this condition is reproduced whenever the number of sites is an integer multiple of four. In more realistic systems, like hydrogen chains, where more orbitals are involved, a richer physics is also present: this includes AFM ordering and metal-to-insulator transitions, as described in Ref. Motta et al. (2020), and also the presence of purely electronic CDW states, which may be related to excitonic mechanisms as described in Ref. Barborini et al. (2022). Within our simple lattice model we manage to verify that a small ratio $V/U$ between local and non-local terms is related to a dominant AFM character, while the CDW character of the model is enhanced at $V/U\simeq 1$.
@@
Finally, we study how broken symmetry solutions are translated in the framework of Kohn-Sham density functional theory (KS-DFT). Here, the evaluation of $G_{KS}$ via the Sham-Schlüter equation, Eq. (II.2), allows one to obtain a non-interacting GF having the same charge density of the interacting system. We verify that the unphysical broken symmetry solutions found by approximate MBPT methods can also be inherited at the KS-DFT level. Moreover, the study of a $N=16$ chain shows that if the model of Eq. (20) has a non-interacting metallic ground state then the derivative discontinuity in the corresponding KS potential tends to be very large, further stressing the difficulties in using the KS energy levels as an approximation to the charged excitation energies of the system.
@@ Appendix A SOP representation of the HF, GW, and 2B self-energies
In this Appendix we report the SOP representation of the self-energy approximations used throughout our work, namely HF, GW, and 2B. Throughout this section, repeated spatial indexes $i$, $j$ and $k$ are meant to be summed over. Since Green’s function and self-energy operators are diagonal in their spin indexes we do not write them explicitly, unless otherwise stated. The Green’s function of the interacting system $G(\omega)$, the RPA polarizability $P(\omega)$, and the screened interaction $W(\omega)$ are defined by the following convention about poles and residues:
$G_{ij}(\omega)$ $=\sum_{s}\frac{A^{s}_{ij}}{\omega-z_{s}}$ (26)
$P_{ij}(\omega)$ $=\sum_{t}\frac{R^{t}_{ij}}{\omega-\Omega_{t}}$ (27)
$W_{ij}(\omega)$ $=\sum_{t}\frac{B^{t}_{ij}}{\omega-w_{t}}$ (28)
where the indexes $i,j$ refers to the matrix elements of the operator in the basis of the lattice coordinates. The density-matrix $\gamma_{ij}$ is related to $G(\omega)$ by:
@@
where the “occupied” label means poles with energies below the Fermi level. The Hartree-Fock self-energy is then given by
@@
where $n_{k}$ is the total charge density on site $k$, while $\gamma_{ij}$ is the spin-resolved density matrix. The $v_{ij}$ are the matrix elements of the electron-electron repulsion term in the basis of the lattice coordinates.
The 2B self-energy is obtained by adding two second order terms to the HF self-energy. The SOP expression of the second order direct term is:
$\Sigma^{2B,d}_{ij}(\omega)=-\int\frac{d\omega_{1}}{2\pi i}e^{i\omega_{1}0^{+}}G_{ij}(\omega+\omega_{1})v_{ik}P_{kl}(\omega_{1})v_{lj}=\sum_{s}^{occ}\sum_{t}^{emp}\frac{A^{s}_{ij}v_{ik}R^{t}_{kl}v_{lj}}{\omega-(z_{s}-\Omega_{t})}-\sum_{s}^{emp}\sum_{t}^{occ}\frac{A^{s}_{ij}v_{ik}R^{t}_{kl}v_{lj}}{\omega-(z_{s}-\Omega_{t})}.$ (31)
Instead, the second order exchange term is given by:
$\Sigma^{2B,x}_{ij}(\omega)$ $=\int\frac{d\omega_{2}}{2\pi i}e^{i\omega_{1}0^{+}}\int\frac{d\omega_{2}}{2\pi i}e^{i\omega_{2}0^{+}}G_{ik}(\omega+\omega_{1})G_{kl}(\omega_{1}+\omega_{2})G_{lj}(\omega_{2})v_{il}v_{kj}$ (32)
@@
In particular from the SOP representation we can directly observe that the direct term of Eq. (31) and the exchange term of Eq. (32) share exactly the same poles. However, the residues of the exchange term are all negative semi-definite, while the residues of the direct term are all positive semi-definite. The positive semi-definiteness of the 2B self energy depends then on the residues of the direct term having a larger trace than the residues of the exchange term. Finally the SOP representation of the GW self-energy is given by:
@@
$=\sum_{s}^{occ}\sum_{t}^{emp}\frac{A^{s}_{ij}B^{t}_{ij}}{\omega-(z_{s}-w_{t})}-\sum_{s}^{emp}\sum_{t}^{occ}\frac{A^{s}_{ij}B^{t}_{ij}}{\omega-(z_{s}-w_{t})}.$
@@ Appendix B SSE in the SOP formalism
$\sum_{j}\int\frac{d\omega}{2\pi i}e^{i\omega 0^{+}}$ $G^{KS}_{ij}(\omega)v^{xc}_{j}G_{ji}(\omega)=$ (34)
$=\sum_{jk}\int\frac{d\omega}{2\pi i}e^{i\omega 0^{+}}G^{KS}_{ij}(\omega)\Sigma_{jk}(\omega)G_{ki}(\omega),$
where the linearized version can be obtained by just substituting $G\rightarrow G_{KS}$ everywhere, including the implicit dependency of the GF of the self-energy. By introducing the following notation for the left-hand-side (LHS) of Eq. (34)
$A_{ij}=\int\frac{d\omega}{2\pi i}e^{i\omega 0^{+}}G^{KS}_{ij}(\omega)G_{ji}(\omega)$ (35)
@@
$b_{i}=\sum_{jk}\int\frac{d\omega}{2\pi i}e^{i\omega 0^{+}}G^{KS}_{ij}(\omega)\Sigma_{jk}(\omega)G_{ki}(\omega),$ (36)
@@
$G^{KS}_{ij}(\omega)$ $=\sum_{s}\frac{P^{s}_{ij}}{\omega-z^{KS}_{s}}$ (38)
$G_{ij}(\omega)$ $=\sum_{s}\frac{R^{s}_{ij}}{\omega-z_{s}}$ (39)
@@
$A^{SSE}_{ij}=\sum_{s}^{emp}\sum_{q}^{occ}\frac{P^{s}_{ij}R^{q}_{ji}}{z_{q}-z^{KS}_{s}}-\sum_{s}^{occ}\sum_{q}^{emp}\frac{P^{s}_{ij}R^{q}_{ji}}{z_{q}-z^{KS}_{s}}.$ (41)
Following Eq. (41), a possible divergence of the matrix elements of $A$, due to $z_{q}=z^{KS}_{s}$ can be avoided if $G$ and $G^{KS}$ have the same quasiparticle HOMO, since this would avoid that “empty” poles of one operator overlap with “occupied” poles of the other. Moreover, this alignment of the chemical potentials also makes the linear system defined in Eq. (37) non rank-defective, since it eliminates the indeterminacy of the potential in terms of a spatially-constant shift.
@@
$b^{SSE}_{i}$ $=$ $\sum_{s}^{emp}\sum_{q}^{occ}\frac{P^{s}_{ij}\Sigma_{jk}(z_{q})R^{q}_{ki}}{z_{q}-z^{KS}_{s}}$
-- $\sum_{s}^{occ}\sum_{q}^{emp}\frac{P^{s}_{ij}\Sigma_{jk}(z^{KS}_{s})R^{q}_{ki}}{z_{q}-z^{KS}_{s}}$
$+$ $\sum_{m}^{occ}G^{KS}_{ij}(\Omega_{m})\Gamma^{m}_{jk}G_{ki}(\Omega_{m})+\sum_{s}^{occ}\sum_{q}^{occ}\tilde{b}^{sq}_{i}.$
@@
We note that the expression of the term $\tilde{b}_{i}^{sq}$ changes according to the presence of a possible pole of second order in the vector $b_{i}$ of Eq. (36). These are obtained in the case of $z_{q}=z^{KS}_{s}$, when both poles are below or at the chemical potential. In this situation the term can be evaluated as
@@
$\tilde{b}_{i}^{sq}=\left\{\begin{array}[]{ll}\frac{P^{s}_{ij}\left[\Sigma_{jk}(z_{q})-\Sigma_{jk}(z^{KS}_{s})\right]R^{q}_{ki}}{z_{q}-z^{KS}_{s}}&\quad\text{if}\quad z_{q}\neq z^{KS}_{s}\\[7.0pt] P^{s}_{ij}\dot{\Sigma}_{jk}(z_{q})R^{q}_{ki}&\quad\text{if}\quad z_{q}=z^{KS}_{s}\end{array}\right.$ (43)

### v2

@@ V Long-range interaction
We note in passing that different parametrizations have also been employed in similar models accounting for a non-local electron-electron repulsion [70, 71, 72]. All these parametrizations, however, share the same $V_{ij}\propto 1/R_{ij}$ behavior at large distances. The model depends on two dimensionless parameters: the ratio $U/t$ defines the strength of the correlation, while the ratio $V/U$ sets the strength of the non-local interaction term with respect to the local one. We take $V/U\leq 1$ as we expect inter-site electron repulsion to be always smaller than the on-site one.
@@ V.1 Results
The nature of the exact ground state can be further characterized by the introduction of appropriate correlation functions involving two density operators. In particular, even if the exact ground state shares the same symmetry of the non-interacting one, it can still develop an AFM or CDW character, which can be measured by the evaluation of ad hoc correlation functions. This is, for instance, the case of the Hubbard model, whose ground state develops strong AFM correlations in the large $U/t$ regime without breaking explicitly the symmetry of the Hamiltonian [7, 8].
@@
which measures the overall charge dimerization on two adjacent sites. The factor of $1/4$ in Eq. (23) takes into account the number of all spin combinations entering the sum and makes Eq. (23) comparable to Eq. (22). The smaller the value of the correlation functions in Eqs. (22) and (23), the larger the corresponding AFM or CDW character. While these correlation functions are easily accessible in the exact calculations, as they can be directly computed through the action of field operators on the many-body wavefunctions, they can not be directly evaluated within the present MBPT approach, where we have mostly access to the one-particle Green’s function ($C^{AFM}$ in Eq. (22) is easily accessible only in the Hubbard model, where the local nature of the electron-electron interaction allows one to extract this correlation function from the knowledge of the total and interaction energies [7, 8]). For the broken symmetry solutions obtained within approximate MBPT calculations, we then consider a “factorized” version of the former correlation functions:
@@
Importantly, for this particular system the symmetry breaking at the MBPT level, while unphysical, follows from the development of analogous correlations in the exact system. Interpreting the ratio $V/U$ in terms of the distance between the sites of the chain – i.e. a small value of $V/U$ representing a large inter-site distance we can see that AFM correlations dominate in the large distance regime as observed for hydrogen chains [1], while CDW correlations are relevant at smaller inter-site distances [3].
@@ V.2 Solutions of the SSE
It is interesting to stress that, due to the simplified nature of the lattice model, the symmetric solution leads to a uniform charge density, which is then the same for all the values of the model parameters. This means that, for the symmetric solutions, the KS potential $v^{\text{KS}}$ is just a constant term that aligns the highest occupied molecular (quasiparticle) orbital (HOMO) of the non-interacting ground state to the HOMO of the interacting symmetric state. In particular, in the case of chains having a metallic non-interacting ground state (see Supplemental Material [42]), the Kohn-Sham Green’s function $G^{\text{KS}}$ describes a metallic system, while the system displays a gap at the MBPT level. We stress here that in order to converge the GW calculations in the case of the symmetric state, the self-consistency loop is initialized with a gapped GF.
Instead, in the case of broken symmetry solutions, $G^{\text{KS}}$ describes a gapped system at finite values of $U$ and $V$. The ratio between the derivative discontinuity $\Delta_{xc}$ and the quasiparticle energy gap is reported in Fig. 8.
@@
We observe that the derivative discontinuity for the $N=16$ chain is large and accounts for $70\%$ to $80\%$ of the quasiparticle energy gap. Similar results have also been observed in model 1D semiconductors with long-range electronic interactions [41]. As an interesting point, the ratio between $\Delta_{xc}$ and $E_{g}^{QP}$ gets smaller by increasing the parameter $V/U$ for the CDW solution, while the opposite happens for the AFM solution. For the $N=16$ chain the value of $V/U$ at which the two curves of Fig. 8 cross is also the value of $V/U$ at which the energy of the AFM solution becomes larger than that of the CDW one at the level of GW. Having a large derivative discontinuity, the energy gap of $G_{KS}$ in the case of broken symmetry solutions is severely underestimated with respect to the quasiparticle gap.
In Fig. 9 we report the DOS of the periodic $N=16$ chain at $U/t=4$ and $V/U=0.50$. The DOS has been evaluated in the case of the GW for the symmetric as well as broken symmetry solutions. The green line reports the DOS of the self-consistent GF, while the red line reports the DOS of $G^{\text{KS}}$. The position of the HOMO level is also highlighted. As discussed in Sec. III.2, the system of linear equations defined by the SSE, Eq. (II.2), is solved requiring that the HOMO level of $G^{\text{KS}}$ is aligned to that of $G$.
Figure 9: DOS of the $N=16$ chain with periodic boundary conditions at $U/t=4$ and $V/U=0.5$ for a long-range Coulomb interaction. We report the DOS for all states with different symmetry. The green line is related to the GW DOS, while the red line is the DOS of the corresponding non-interacting $G^{\text{KS}}$. The HOMO quasiparticle level is marked with a downwards triangle and shifted to $\omega=0$ in all cases. The frequency is scaled with the hopping parameter $t$ and the DOS is reported in arbitrary units.
@@ VI Conclusions
In this work we study one-dimensional lattice models with an approximate many-body perturbation theory (MBPT) treatment of the electron-electron interaction. The Dyson equation is solved in a fully self-consistent manner via the algorithmic-inversion method based on the sum-over-pole representation of dynamical operators (AIM-SOP) [24, 25, 26]. Our calculations demonstrate that the AIM-SOP framework enables to solve self-consistently the Dyson equation in a stable and accurate way, for the system properties considered, obtaining an excellent agreement with the existing literature [28, 29, 7]. As a result, the fully self-consistent implementation allows one to study in detail the total energy of broken symmetry solutions by evaluating unambiguously the value of the Klein functional at self-consistency. Notably, in this work we make use of – and we further numerically validate – a recent analytical expression to compute the term $\text{Tr}_{\omega}\text{Ln}\{G_{0}^{-1}G\}$ appearing in the Klein functional [40].
@@
We then apply the AIM-SOP framework to study broken symmetry solutions, extending the results reported in Refs. [7, 8], to model systems accounting also for non-local electronic interactions. We show that – while the exact ground state has a homogeneous charge density as in the case of the Hubbard model – broken symmetry solutions with a charge density wave (CDW) or antiferromagnetic (AFM) ordering can be found within an approximate MBPT treatment. We consider the example of a 4-site chain where the unphysical MBPT broken symmetry solutions can be traced back to the development of an AFM or CDW character in the exact ground state, which is related to the interplay between the local and the non-local electronic interactions and can be probed by density-density correlation functions such as those of Eq. (22) and Eq. (23).
In particular, the CDW character gets enhanced when the strength of the non-local electronic interaction $V$ approaches the strength of the local electronic repulsion $U$. For the 4-site chain we show that the ground state switches from AFM to CDW when the ratio between the local and non-local interaction of Eq. (21) becomes larger than $V/U=0.75$, independently of the value of $U/t$. Moreover, the estimation of the total energy with the GW approximation improves considerably for symmetry unrestricted calculations, both for AFM and CDW solutions. We also observe that broken symmetry solutions in longer chains are found when the non-interacting ground state of the model Hamiltonian in Eq. (20) is metallic.
@@
Finally, we study how the broken symmetry solutions are translated in the framework of Kohn-Sham density functional theory (KS-DFT) by adopting the AIM-SOP framework to solve the (non-linearized) Sham-Schlüter equation. We verify that in this model, if the non-interacting ground state is metallic, the KS solution is always metallic in case of symmetry restricted calculations, while a small gap opens at the KS level of theory for broken symmetry solutions. Furthermore, we show for a 16-site chain that if the model of Eq. (20) has a non-interacting metallic ground state, then the derivative discontinuity in the corresponding KS potential tends to be remarkably large, further stressing the relevance of this exchange-correlation driven correction to the KS gap.
@@ Appendix A SOP representation of the HF, GW, and 2B self-energies
In this Appendix we report the SOP representation of the self-energy approximations used throughout our work, namely HF, GW, and 2B. Throughout this section, repeated indexes $i$, $j$ and $k$ are meant to be summed over. Since Green’s function and self-energy operators are diagonal in their spin indexes we do not write them explicitly, unless otherwise stated. The Green’s function of the interacting system $G(\omega)$, the RPA polarizability $P(\omega)$, and the screened interaction $W(\omega)$ are defined by the following convention about poles and residues:
$G_{ij}(\omega)$ $=\sum_{s}\frac{A^{s}_{ij}}{\omega-z_{s}},$ (26)
$P_{ij}(\omega)$ $=\sum_{t}\frac{R^{t}_{ij}}{\omega-\Omega_{t}},$ (27)
$W_{ij}(\omega)$ $=\sum_{t}\frac{B^{t}_{ij}}{\omega-w_{t}},$ (28)
where the indexes $i,j$ refers to the lattice coordinates. The density-matrix $\gamma_{ij}$ is related to $G(\omega)$ by:
@@
where the “occupied” label indicates poles with energies below the Fermi level. The Hartree-Fock self-energy is then given by
@@
where $n_{k}$ is the total charge density on site $k$, $\gamma_{ij}$ is the spin-resolved density matrix, and $v_{ij}$ are the matrix elements of the electron-electron repulsion term in the basis of the lattice coordinates.
The 2B self-energy is obtained by adding two second order terms (direct and exchange) to the HF one.
$\Sigma^{2B,d}_{ij}(\omega)$ $=$ $-\int\frac{d\omega_{1}}{2\pi i}G_{ij}(\omega+\omega_{1})\,v_{ik}\,P_{kl}(\omega_{1})v_{lj}=\sum_{s}^{occ}\sum_{t}^{emp}\frac{A^{s}_{ij}v_{ik}R^{t}_{kl}v_{lj}}{\omega-(z_{s}-\Omega_{t})}-\sum_{s}^{emp}\sum_{t}^{occ}\frac{A^{s}_{ij}v_{ik}R^{t}_{kl}v_{lj}}{\omega-(z_{s}-\Omega_{t})}.$ (31)
$\Sigma^{2B,x}_{ij}(\omega)$ $=$ $\phantom{-}\int\frac{d\omega_{1}}{2\pi i}\int\frac{d\omega_{2}}{2\pi i}G_{ik}(\omega+\omega_{1})\,G_{kl}(\omega_{1}+\omega_{2})\,G_{lj}(\omega_{2})\,v_{il}\,v_{kj}$
@@
In particular from the SOP representation we can directly observe that the direct term of Eq. (31) and the exchange term of Eq. (A) share exactly the same poles. However, the residues of the exchange term are all negative semi-definite, while the residues of the direct term are all positive semi-definite. The positive semi-definiteness of the 2B self energy depends then on the residues of the direct term being (matrix-wise) larger than the residues of the exchange term. Finally the SOP representation of the GW self-energy is given by:
@@
$=$ $\phantom{-}\sum_{s}^{occ}\sum_{t}^{emp}\frac{A^{s}_{ij}B^{t}_{ij}}{\omega-(z_{s}-w_{t})}$
$-\sum_{s}^{emp}\sum_{t}^{occ}\frac{A^{s}_{ij}B^{t}_{ij}}{\omega-(z_{s}-w_{t})}.$
@@ Appendix B SSE in the SOP formalism
$\sum_{j}\int\frac{d\omega}{2\pi i}e^{i\omega 0^{+}}G^{\text{KS}}_{ij}(\omega)\,v^{\text{KS},xc}_{j}\,G_{ji}(\omega)=\\ =\sum_{jk}\int\frac{d\omega}{2\pi i}e^{i\omega 0^{+}}G^{KS}_{ij}(\omega)\Sigma_{jk}(\omega)G_{ki}(\omega),$ (34)
where the linearized version can be obtained by just substituting $G\rightarrow G^{\text{KS}}$ everywhere, including the implicit dependency of the GF of the self-energy. By introducing the following notation for the left-hand-side (LHS) of Eq. (34)
$A_{ij}=\int\frac{d\omega}{2\pi i}e^{i\omega 0^{+}}\,G^{\text{KS}}_{ij}(\omega)\,G_{ji}(\omega)$ (35)
@@
$b_{i}=\sum_{jk}\int\frac{d\omega}{2\pi i}e^{i\omega 0^{+}}\,G^{\text{KS}}_{ij}(\omega)\,\Sigma_{jk}(\omega)\,G_{ki}(\omega),$ (36)
@@
$G^{\text{KS}}_{ij}(\omega)$ $=\sum_{s}\frac{P^{s}_{ij}}{\omega-z^{\text{KS}}_{s}},$ (38)
$G_{ij}(\omega)$ $=\sum_{s}\frac{R^{s}_{ij}}{\omega-z_{s}},$ (39)
@@
$A_{ij}=\sum_{s}^{emp}\sum_{q}^{occ}\frac{P^{s}_{ij}R^{q}_{ji}}{z_{q}-z^{\text{KS}}_{s}}-\sum_{s}^{occ}\sum_{q}^{emp}\frac{P^{s}_{ij}R^{q}_{ji}}{z_{q}-z^{\text{KS}}_{s}}.$ (41)
Following Eq. (41), a possible divergence of the matrix elements of $A$, due to $z_{q}=z^{\text{KS}}_{s}$ can be avoided if $G$ and $G^{\text{KS}}$ have the same quasiparticle HOMO, since this would avoid that “empty” poles of one operator overlap with “occupied” poles of the other. Moreover, this alignment of the chemical potentials also makes the linear system defined in Eq. (37) non rank-defective, since it eliminates the indeterminacy of the potential in terms of a spatially-constant shift.
@@
$b_{i}$ $=$ $\sum_{s}^{emp}\sum_{q}^{occ}\frac{P^{s}_{ij}\Sigma_{jk}(z_{q})R^{q}_{ki}}{z_{q}-z^{\text{KS}}_{s}}$
-- $\sum_{s}^{occ}\sum_{q}^{emp}\frac{P^{s}_{ij}\Sigma_{jk}(z^{\text{KS}}_{s})R^{q}_{ki}}{z_{q}-z^{\text{KS}}_{s}}$
$+$ $\sum_{m}^{occ}G^{\text{KS}}_{ij}(\Omega_{m})\,\Gamma^{m}_{jk}\,G_{ki}(\Omega_{m})+\sum_{s}^{occ}\sum_{q}^{occ}\tilde{b}^{sq}_{i}.$
@@
We note that the expression of the term $\tilde{b}_{i}^{sq}$ changes according to the presence of a possible pole of second order in the vector $b_{i}$ of Eq. (36). These are obtained in the case of $z_{q}=z^{\text{KS}}_{s}$, when both poles are below or at the chemical potential. In this situation the term can be evaluated as
@@
$\tilde{b}_{i}^{sq}=\left\{\begin{array}[]{ll}\frac{P^{s}_{ij}\,\left[\Sigma_{jk}(z_{q})-\Sigma_{jk}(z^{\text{KS}}_{s})\right]\,R^{q}_{ki}}{z_{q}-z^{\text{KS}}_{s}}&\quad\text{if}\quad z_{q}\neq z^{\text{KS}}_{s}\\[7.0pt] P^{s}_{ij}\,\dot{\Sigma}_{jk}(z_{q})\,R^{q}_{ki}&\quad\text{if}\quad z_{q}=z^{\text{KS}}_{s}\end{array}\right.$ (43)